OpenStudy (mtimko):

How long (in s) does it take for a radio signal from the earth to reach an astronaut on the moon? The distance from the earth to the moon is 2.4 x 105 mi.

OpenStudy (arindameducationusc):

t=s/c where t= time taken (we need to find) s=2.4*105 miles (convert to meters) c=speed of light(3*10^8)

OpenStudy (mtimko):

How did you know to use speed of light?

OpenStudy (arindameducationusc):

radio signals are Electromagnetic waves, Even light is Electromagnetic waves. You will get your answer.....Sure.... @mtimko

OpenStudy (arindameducationusc):

If it was useful, don't forget to medal.... and fan....